Assume a firm has a market value of $500, it has 50 units of capital with a price of $8 per

unit. Compute Tobin’s q to determine whether the firm should invest in more capital.

q = __________. Invest? (Yes or no) __________
Answer #1

Cost of replacement of capital = 50*8 = $400

Tobin's q = Market value/Cost of replacement of capital = 500/400 = 1.25
So, it should invest as q > 1.

q = 1.25
Invest - Yes